OpenStudy (anonymous):

Determine f(2), f(t-4). a.) f(x) = x^2 + 9 b.) f(x) = x^2 + 3x + 2

OpenStudy (anonymous):

plz explain

OpenStudy (anonymous):

\[f(\color{red}{x}) =\color{red}{x^2}+9\\f(\color{red}{2}) =\color{red}{2^2}+9=13\] can you do the rest?

OpenStudy (anonymous):

that yes but the t-4 part no

OpenStudy (tylerd):

plug in t-4 for x

OpenStudy (tylerd):

see what you get

OpenStudy (xapproachesinfinity):

\(\huge \color{green}{f(\color{red}{t-4})=[~~~~~~]^2+9}\) in the barackets put t-4 instead of x \(\huge \color{green}{f(\color{red}{t-4})=[\color{red}{t-4}]^2+9}\)

OpenStudy (xapproachesinfinity):

can you do it now?

OpenStudy (xapproachesinfinity):

instead of x you put t-4, the same if it is 2 or any other number or expression

OpenStudy (anonymous):

but wat is (t-4)^2

OpenStudy (anonymous):

im not suppose to use a calculator

OpenStudy (xapproachesinfinity):

you don't know how to expand that \(\huge \color{brown}{(t-4)^2=(t-4)(t-4)}\) Foil and you have \(\huge \color{brown}{(t-4)(t-4)=t^2-4t-4t-4(-4)\\~~~~~~~~~~~~~~~~~~~~~=t^2-8t+16}\)

OpenStudy (xapproachesinfinity):

Do you know how to do this?

OpenStudy (xapproachesinfinity):

\(\huge \color{brown}{(a+b)^2=a^2+2ab+b^2}\) this is called binomial expansion you should be able to this by now since you are taking that course!

OpenStudy (anonymous):

ok i got ti now thank you

OpenStudy (xapproachesinfinity):

\(\Huge \frak\color{blueviolet}{\bigstar~you~are ~welcome\bigstar}\)